# High-order Computation of Floquet Multipliers and Subspaces using Multistep Methods

### v1

@@ 3.1 Derivation of the Periodic-PEP
$\operatorname{col}(x_{i})_{i=m+1}^{m+l}=\left[x_{m+1}^{\mathrm{H}},\,x_{m+2}^{\mathrm{H}},\,\ldots,\,x_{m+l}^{\mathrm{H}}\right]^{\mathrm{H}}\in\mathbb{R}^{ln},$ (17)
which collects lm+1l-m+1 consecutive state vectors ending at step $i$. We then define the block-diagonal matrices $\mathcal{A}_{j}=\operatorname{diag}_{i}(A_{j}^{(i)}),\,j=0,\ldots,d-1,$ and the cyclic shift operator $\mathcal{C}=C_{p}\otimes I_{n}$, $C_{p}=[e_{2},\dots,e_{p},e_{1}],$ where $e_{i}$ denotes the $i$-th standard basis vector of $\mathbb{R}^{p}$.
With these notations, the periodic-PEP in definition˜4 can be equivalently written in the compact matrix form.
$\left(\theta^{d}I_{np}-\sum_{j=0}^{d-1}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)\operatorname{col}(u^{(i)})_{i=1}^{p}=0.$ (18)
3.2 Equivalent Form of the Periodic-PEP using Discrete State Transition Matrix
This subsection transforms the periodic-PEP in eq.˜18 into an equivalent form to reveal its connection with the general solution of the underlying difference equation. We show that through a suitable linearization and a structured reordering of the companion-form matrices, eq.˜18 becomes an $nd$-dimensional periodic eigenvalue problem. The coefficient matrices of this new problem are identical to the state transition matrices of the difference equation eq.˜13. This identity establishes a direct correspondence between the eigenpairs of the periodic-PEP and the fundamental solutions of the periodic difference system.
The companion form of the PEP eq.˜18 is:
$\begin{bmatrix}&I_{np}&&\\ &&\ddots&\\ &&&I_{np}\\ \mathcal{A}_{0}\mathcal{C}^{d}&\mathcal{A}_{1}\mathcal{C}^{d-1}&\cdots&\mathcal{A}_{d-1}\mathcal{C}\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}\theta,$ (19)
The eigenvector in eq.˜19 exhibits a two-level structure: the inner structure encodes the periodicity of the coefficients, while the outer structure arises from the linearization of eq.˜18. To make this structure explicit, we introduce a permutation matrix $\mathcal{P}$ that reorders the stacked vector:
After this reordering, the eigenvalue problem takes the following block-cyclic form:
$\begin{bmatrix}&&&L^{(1)}\\ L^{(2)}&&&\\ &\ddots&&\\ &&L^{(p)}&\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}\theta,$ (20)
where each block $L^{(i)}$ is the companion matrix
$L^{(i)}=\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}$. (21)
The block-cyclic structure implies that, for every $i=1,\ldots,p$, the eigenvalue problem reduces to the $nd$-dimensional periodic eigenvalue problem:
$L^{(i)}L^{(i-1)}\cdots L^{(i-p+1)}\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}\theta^{p}.$ (22)
This transformation maps each $p$-fold set of eigenvalues of definition˜4 to a single eigenvalue $\theta^{p}$, effectively removing the rotational symmetry and revealing the essential dynamics.
Recall eq.˜12, we observe that each companion matrix $L^{(i)}$ describes one-step propagation of the stacked state vectors:
$\operatorname{col}(x_{j})_{j=i-d+1}^{i}=L^{(i)}\operatorname{col}(x_{j})_{j=i-d}^{i-1}.$ (23)
To describe the evolution, we define the discrete STM in a $dn$-dimensional state space [21]:
$F_{i+j,i}=L^{(i+j)}L^{(i+j-1)}\cdots L^{(i+1)}.$ (24)
Because the coefficients $L^{(i)}$ are $p$-periodic, we have $F_{i+kp,i+(k-1)p}=F_{i+p,i}$ for all integers $k$. Therefore, we partition the stacked sequence $\operatorname{col}(x_{j})_{j=i-d+1}^{i}$ for all $i\in\mathbb{Z}$ into $p$ groups according to the remainder of the index $i$ modulo $p$. Each group forms a phase slice $\operatorname{col}(x_{j})_{j=i+kp-d+1}^{i+kp}$, and within each slice the evolution satisfies a constant-coefficient difference equation:
$\operatorname{col}(x_{j})_{j=i+kp-d+1}^{i+kp}=F_{i+p,i}\,\operatorname{col}(x_{j})_{j=i+(k-1)p-d+1}^{i+(k-1)p}.$ (25)
Here, $F_{i+p,i}$ acts as the discrete state-transition matrix over one period for the corresponding slice, and the associated eigenvalue problem is precisely eq.˜22.
Moreover, these phase slices are not independent, because the transition matrices for different starting points are all similar:
$F_{i+p,i}=F_{i,0}F_{p,0}F_{i,0}^{-1},$ (26)
and therefore share the same spectrum. As a result, the eigen-decomposition of $F_{p,0}$ and state transition $F_{i,0}$ fully determines the eigen-structure of all $F_{i+p,i}$.
In this section, we have shown the discrete STM over a period $F_{i+p,i}$ reveals structural properties that closely mirror those of the continuous STM $\Phi(T,0)$ discussed in section˜2.1. The key distinction is that the discrete STM acts on an $nd$-dimensional stacked state corresponding to the difference equation, whereas the continuous STM evolves in the original $n$-dimensional state space of the LPTV system.
When the coefficient matrix $A_{j}^{(i)}$ are from multistep discretization of the LPTV system, we expect $n$ of the eigenvalues of $F_{p,0}$ to approximate the $n$ Floquet multipliers of the continuous system. However, the remaining n(d1)n(d-1) eigenvalues of $F_{p,0}$ require further explanation, which will be addressed by introducing spurious systems in the following section.
4 Spurious Systems from Multistep Methods
The continuous LPTV system eq.˜1 evolves in an $n$-dimensional state space. Discretizing it with a $d$-step multistep method yields a difference method whose state space is $nd$-dimensional and one-period STM is $F_{p,0}$. In this section, based on the given LPTV system and the adopted $d$-step multistep method, we design d1d-1 additional $n$-dimensional spurious systems together with their corresponding multistep discretizations. The discretizations of the physical and spurious systems share identical coefficient matrices, hence yield the same difference equation in coefficient form, but differ in the way the state variables enter the recursion. Utilizing the Floquet multipliers and eigenvectors of these spurious systems, we construct an $nd$-dimensional auxiliary STM $\Psi(t_{p},t_{0})$ that serves as the limiting counterpart of $F_{p,0}$ as $\Delta t\to 0$.
4.1 Spurious Systems from Discretizing Scalar System
To illustrate how multistep methods introduce the spurious systems, we first analyze a simple scalar system, $\dot{y}=0$, $t\in[0,T]$. Applying the $d$-step discretization on a periodic time grid leads to a high-order difference equation:
$y_{i}=-\sum_{j=0}^{d-1}\alpha_{j}^{(i)}y_{i-d+j},\quad i=1,\ldots,p.$ (27)
It corresponds to the following periodic-PEP:
$\left(\nu^{d}I_{p}-\sum_{j=0}^{d-1}\mathbf{A}_{j}C^{d-j}\nu^{j}\right)\eta=0,$ (28)
where $\mathbf{A}_{j}=-\operatorname{diag}\left\{\alpha_{j}^{(i)}\right\}_{i=1}^{p}$, and $\eta=[\eta^{(1)},\ldots,\eta^{(p)}]^{\mathrm{H}}$. According to definition˜4, the $d$ periodic eigenvalues of eq.˜28 are denoted by $\nu_{1}^{p},\ldots,\nu_{d}^{p}$. According to definition˜3, the trivial eigenvalue $\nu_{1}=1$ remains unchanged, while all other eigenvalues satisfy $\lvert\nu_{\tau}\rvert<1$ for $\tau=2,\ldots,d$. The corresponding eigenvector of $\nu_{1}$ is still $1,1,\ldots,1]^{\mathrm{H}}$, whereas the other eigenvectors are no longer uniform when the stepsize is not constant.
4.2 Spurious Systems from Discretizing LPTV Systems
Returning now to the general system eq.˜13, each coefficient matrix in eq.˜14 is a small perturbation of $\mathbf{A}_{j}\otimes I_{n}$ when $\Delta t$ is sufficiently small. Motivated by the scalar case, we assume that the general solution of eq.˜13 inherits the same structure, taking the form $x_{\tau,i}\eta^{(i)}\nu^{i}$. Substituting this back to eq.˜13 yields:
$x_{\tau,i}\eta_{\tau}^{(i)}\nu_{\tau}^{i}-\sum_{j=1}^{d}A_{j}^{(i)}x_{\tau,i-d+j}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{i-d+j}=0,\quad\tau=1,\ldots,d.$ (29)
The mode $\nu_{1}=1$ with $\eta_{1}^{(i)}\equiv 1$ corresponds to the physical dynamics and recovers eq.˜13. The remaining modes $\nu_{\tau}$ $(\tau>1)$ also describe valid dynamical components through $x_{\tau,i}$; however, their associated decaying envelopes render them spurious in the context of the original system.
To formalize how $x_{\tau,i}$ captures these spurious dynamics, we associate each mode $\tau$ with a corresponding spurious LPTV system and a spurious multistep method. In doing so, we show that eq.˜29 can be obtained from the discretization of these spurious systems. This motivates the construction of the corresponding spurious multistep methods, whose coefficients are then defined as
$\alpha_{\tau,j}^{(i)}=\frac{\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d}\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}h_{j}^{(i)}},\quad\beta_{\tau,j}=\frac{\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d}\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}.$ (30)
where the envelope terms are incorporated into the numerators, and the normalization ensures the first-order consistency condition (definition˜2). The scaling factor $h_{j}^{(i)}$ in the denominator of $\alpha_{\tau,j}^{(i)}$ is defined by
$h_{j}^{(i)}\equiv\frac{t_{i+j}-t_{i}}{t_{i+d}-t_{i+d-1}}=\frac{\Delta t^{(i+1)}}{\Delta t^{(i+d)}}\sum_{l=1}^{j}\frac{\Delta t^{(i+l)}}{\Delta t^{(i+1)}}=\left(\prod_{l=1}^{d}{\omega^{(i+l)}}^{-1}\right)\sum_{l=1}^{j}\omega^{(i+l)}\cdots\omega^{(i+1)},$ (31)
where $\omega^{(i)}=\Delta t^{(i)}/\Delta t^{(i-1)}$ denotes the local stepsize ratio. As seen from eq.˜31, $h_{j}^{(i)}$ depends only on the stepsize ratios $\omega^{(i+l)}$ for $l=1,\ldots,d$, and reduces to $h_{j}^{(i)}=j$ under constant stepsizes.
We now construct the corresponding spurious LPTV systems:
$\frac{\mathrm{d}x_{\tau}}{\mathrm{d}t}$ $=\rho_{\tau}(t)G(t)x_{\tau}(t),$ (32)
$\rho_{\tau}(t)$ $=\frac{\sum_{j=0}^{d}\beta_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}}{\sum_{j=0}^{d-1}\alpha_{j}^{(i)}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}h_{j}^{(i)}},\quad t\in[t_{i-1},t_{i}],\,\tau=2,\ldots,d.$
Although the underlying system is locally defined on each interval $t_{i-1},t_{i}]$ with potentially distinct dynamics, the $d$-step multistep discretization requires evaluating the equation over the combined interval $t_{i-d},t_{i}]$. To this end, we formally extend the equation form from the last interval $t_{i-1},t_{i}]$ to the preceding intervals $t_{i-d},t_{i-1}]$, effectively using the definition on $t_{i-1},t_{i}]$ to cover the earlier steps. Applying the spurious multistep coefficients from eq.˜30 to this extended system then yields
$\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}I_{n}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)x_{\tau,i-d+j}\eta_{\tau}^{(i-d+j)}\nu_{\tau}^{j}=0,\quad\tau=1,\ldots,d.$ (33)
Left-multiplying by $\left(I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}$ then recovers eq.˜29, and we can write eq.˜23 in an extended version
$L^{(i)}\operatorname{col}\left(x_{\tau,j}\,\eta^{(j)}\,\nu^{\,j}\right)_{j=i-d}^{i-1}=\operatorname{col}\left(x_{\tau,j}\,\eta^{(j)}\,\nu^{\,j}\right)_{j=i-d+1}^{i}.$ (34)
At last, we estimate the local truncation error of the spurious multistep method. Accroding to eq.˜10, we have the following lemma.
Lemma 3.
Consider the spurious $d$-step discretization eq.˜33 applied to the spurious system. When the stepsize is constant, the local truncation error of eq.˜33 is second-order, but its local truncation error drops to first order when stepsize is variable.
4.3 Construction of Auxiliary State Transition Matrix
Similar to the structure of $F_{i+1,i}=L^{(i)}$, we now consider a continuous representation. Let $X_{\tau}(t)$ denote a fundamental solution matrix of eq.˜32 for any fixed full rank initial matrix condition, and we define an $n$-dimensional subspace of an $nd$-dimensional space spanned by:
$\mathcal{X}_{\tau}\equiv\operatorname{span}\left\{\operatorname{col}\left(X_{\tau}^{(i)}(t_{j})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}\right\}.$
Next, we show that the subspaces $\mathcal{X}_{1},\ldots,\mathcal{X}_{d}$ form a direct sum spanning the entire $nd$-dimensional space, that is $\mathcal{X}_{1}\oplus\mathcal{X}_{2}\oplus\cdots\oplus\mathcal{X}_{d}=\mathbb{R}^{(nd)\times(nd)}.$ To see this, we collect the bases of all subspaces into the block matrix
$\begin{bmatrix}X_{1}(t_{i-d+1})&X_{2}(t_{i-d+1})\eta_{2}^{(i-d+1)}&\cdots&X_{d}(t_{i-d+1})\eta_{d}^{(i-d+1)}\\ \vdots&\vdots&\ddots&\vdots\\ X_{1}(t_{i-1})&X_{2}(t_{i-1})\eta_{2}^{(i-1)}\nu_{2}^{i-1}&\cdots&X_{d}(t_{i-1})\eta_{d}^{(i-1)}\nu_{d}^{i-1}\\ X_{1}(t_{i})&X_{2}(t_{i})\eta_{2}^{(i)}\nu_{2}^{i}&\cdots&X_{d}(t_{i})\eta_{d}^{(i)}\nu_{d}^{i}\\ \end{bmatrix}$. (35)
Assume that $X_{\tau}(t_{i})$ is nonsingular for $\tau=1,\ldots,d$. For a fixed $\tau$ and sufficiently small $\Delta t$, the matrices $X_{\tau}(t_{j})$ with $j=i-d+1,\ldots,i$ differ from $X_{\tau}(t_{i})$ only by $\mathcal{O}(\Delta t)$, and the scaling factor $\eta^{(j)}$ vary smoothly as well. Consequently, the assembled block matrix can be regarded as a small perturbation of a block Vander monde matrix, which remains nonsingular provided that $\nu_{\tau}$ are pairwise distinct.
We define an auxiliary STM $\Psi(t_{i},t_{i-1})$ that satisfies
$\operatorname{col}\left(x_{\tau}^{\mathrm{H}}(t_{j})\,\eta_{\tau}^{(j)}\,\nu_{\tau}^{\,j}\right)_{j=i-d+1}^{i}=\Psi(t_{i},t_{i-1})\operatorname{col}\left(x_{\tau}^{\mathrm{H}}(t_{j})\,\eta_{\tau}^{(j)}\,\nu_{\tau}^{\,j}\right)_{j=i-d}^{i-1}.$ (36)
The block companion form of $\Psi(t_{i},t_{i-1})$ on subspace $\mathcal{X}_{\tau}$ is given by
$\left.\Psi(t_{i},t_{i-1})\right|_{\mathcal{X}_{\tau}}=\begin{bmatrix}0&I_{n}&&\\ &\ddots&\ddots&\\ &&0&I_{n}\\ 0&\cdots&0&\frac{\eta_{\tau}^{(i)}}{\eta_{\tau}^{(i-1)}}\nu_{\tau}\Phi_{\tau}(t_{i},t_{i-1})\\ \end{bmatrix}$. (37)
Repeated transition over one full period yields
$\left.\Psi(t_{i+p},t_{i})\right|_{\mathcal{X}_{\tau}}=$ $\begin{bmatrix}0&\cdots&0&\frac{\eta_{\tau}^{(i-d+1)}}{\eta_{\tau}^{(i)}}\Phi_{\tau}(t_{i+p-d+1},t_{i})\nu_{\tau}^{1-d}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&0&\frac{\eta_{\tau}^{(i-1)}}{\eta_{\tau}^{(i)}}\Phi_{\tau}(t_{i+p-1},t_{i})\nu_{\tau}^{-1}\\ 0&\cdots&0&\Phi_{\tau}(t_{i+p},t_{i})\\ \end{bmatrix}$. (38)
Let $U_{\tau}(t_{i})$ be the eigenvector matrix of the STM of the $\tau$-th system $\Phi_{\tau}(t_{i+p},t_{i})$ associated with the Floquet multipliers $e^{M_{\tau}T}$ satisfying $\Phi_{\tau}(t_{i+p},t_{i})U_{\tau}(t_{i})=U_{\tau}(t_{i})e^{M_{\tau}T}.$ Denote the stacked eigenvector matrix:
$\mathcal{U}_{\tau}(t_{i})=\operatorname{col}\left(\Phi_{\tau}(t_{j},t_{i})U_{\tau}(t_{i})\eta_{\tau}^{(j)}\nu_{\tau}^{j}\right)_{j=i-d+1}^{i}.$ (39)
Then, it follows that $\mathcal{U}_{\tau}(t_{i+p})=\mathcal{U_{\tau}}(t_{i})\nu_{\tau}^{p}e^{M_{\tau}T}$, and
$\Psi(t_{i+p},t_{i})\mathcal{U_{\tau}}(t_{i})=\mathcal{U_{\tau}}(t_{i})\nu_{\tau}^{p}e^{M_{\tau}T}.$ (40)
The case $\tau=1$ corresponds to $\nu_{\tau}=1$, which corresponds to the physical system, whose eigenvalues $e^{M_{1}T}$ are the Floquet multipliers of the original LPTV system eq.˜1. The remaining $\tau=2,\ldots,d$ correspond to the spurious systems, whose eigenvalues $\nu_{\tau}^{p}e^{M_{\tau}T}$ are scaled versions of the Floquet multipliers of the spurious LPTV systems eq.˜32, and thus decay geometrically with increasing $p$ since $\lvert\nu_{\tau}\rvert<1$.
5 Spectral Convergence Analysis of Periodic-PEP from Multistep Discretization
To establish the convergence behavior of the discrete STM $F_{p,0}$ towards $\Psi(t_{p},t_{0})$, we analyze the difference $\Psi(t_{p},t_{0})-F_{p,0}$ acting on $\mathcal{U}_{\tau}$ for physical system $\tau=1$ and spurious systems $\tau=2,\ldots,d$. Since all spurious subspaces behave similarly, it suffices to present the proof for the representative two-step case ($d=2$). The proof proceeds in three steps:
  1. 1.
    Estimate the upper bound of $\left\lVert F_{i+l,i}\right\rVert$.
  2. 2.
    Estimate $\left\lVert[\Psi(t_{p},t_{0})-F_{p,0}][\mathcal{U}_{1},\mathcal{U}_{2}]\right\rVert$ by decomposing it into a sum of local differences across steps and bound each term.
  3. 3.
    Express the difference $\left\lVert[\Psi(t_{p},t_{0})-F_{p,0}][\mathcal{U}_{1},\mathcal{U}_{2}]\right\rVert$ in an orthonormal basis aligned with $\Psi(t_{p},t_{0})$’s invariant subspaces and apply the subspace perturbation theorem to conclude convergence of the physical and spurious subspaces.
Before we start the convergence analysis, we first introduce the following assumptions for the LPTV systems, periodic time grid and the multistep methods.
  1. A1.
    Consistency and stability of the multistep method: The $d$-step multistep method is of consistency order $s$ and zero-stable. Its coefficients $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$ are sufficiently smooth functions of the step-size ratios $\omega^{(i)}$.
  2. A2.
    Bounded variable step sizes: The stepsize ratios $\omega^{(i)}$ of the periodic time grid satisfy uniform bounds, i.e., $\omega<\omega^{(i)}<\Omega$, for some constants $0<\omega<1<\Omega$. Consequently, the coefficients $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$ admit uniform upper and lower bounds
    $\alpha_{\min}\leq\alpha_{j}^{(i)}\leq\alpha_{\max},\qquad\beta_{\min}\leq\beta_{j}^{(i)}\leq\beta_{\max},$
    for all $j=0,1,\ldots,d$ and $i=1,\ldots,p$.
  3. A3.
    Smoothness and boundedness of the LPTV system: The periodic matrix $G(t)$ in eq.˜1 is $s$-times continuously differentiable on $[0,T]$. In particular, smoothness implies that both $G(t)$ and the state-transition matrix $\Phi(t,s)$ are uniformly bounded:
    $\|G(t)\|_{\mathrm{F}}\leq C_{G},\qquad\|\Phi(t,s)\|\leq C_{\Phi},\quad\forall\,t-s\in[0,T].$
We begin by establishing the boundedness of the discrete STM $F_{i+p,i}$.
Lemma 4 (Stability and Boundedness of Discrete STM).
Under Assumption A1-A3, let $F_{i+l,i}$ denote the discrete STM over $l$ steps for the multistep discretization of eq.˜1. Then there exists a constant $C_{F}>0$, independent of $l$, such that
$\|F_{i+l,i}\|_{\mathrm{F}}\leq C_{F},\qquad l=1,\ldots,p.$
Proof.
Let $\mathbf{F}_{i+l,i}$ denote the discrete STM associated with the zero scalar system $\dot{y}=0$. By definition˜3, it satisfies $\|\mathbf{F}_{i+l,i}\|_{\mathrm{F}}\leq C_{0}$ for all $l\in\mathbb{N}$. Strictly speaking, the discrete STM acting on the $nd$-dimensional stacked state is $\mathbf{F}_{i+l,i}\otimes I_{n}$; however, for notational simplicity, we continue to denote it by $\mathbf{F}_{i+l,i}$.
We now regard $F_{i+l,i}$ as a perturbation of $\mathbf{F}_{i+l,i}$. For each step,
$\left\lVert F_{i+1,i}-\mathbf{F}_{i+1,i}\right\rVert_{\mathrm{F}}$ $\leq\sum_{j=0}^{d-1}\left\lVert\Delta A_{j}^{(i)}\right\rVert_{\mathrm{F}},$
where $\Delta A_{j}^{(i)}$ for $j=0,\ldots,d-1$ are defined as
$\Delta A_{j}^{(i)}$ $\equiv\left(\alpha_{d}^{(i)}I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}\left(\alpha_{j}^{(i)}I_{n}-\Delta t^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\right)-\alpha_{d}^{-1}\alpha_{j}I_{n}$
$=\Delta t^{(i)}\left(\alpha_{d}^{(i)}I_{n}-\Delta t^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}{\alpha_{d}^{(i)}}^{-1}\left(\beta_{d}^{(i)}\alpha_{j}^{(i)}G^{(i)}-\alpha_{d}^{(i)}\beta_{j}^{(i)}G^{(i+j)}\right).$
For stepsize $\Delta t^{(i)}<\frac{\left\lvert\alpha_{d}^{(i)}\right\rvert}{2\left\lvert\beta_{d}^{(i)}\right\rvert C_{G}}\leq\frac{\alpha_{\max}}{2\beta_{\min}C_{G}}$, we have
$\left\lVert\Delta A_{j}^{(i)}\right\rVert_{\mathrm{F}}$ $<\left\lvert\alpha_{d}^{(i)}\right\rvert^{-1}\frac{2\alpha^{(i)}\beta^{(i)}C_{G}\Delta t^{(i)}}{\left\lvert\alpha_{d}^{(i)}\right\rvert-\Delta t\left\lvert\beta_{d}^{(i)}\right\rvert C_{G}}\leq 4\alpha_{\min}^{-2}\alpha_{\max}\beta_{\max}C_{G}\Delta t^{(i)},$
Hence, the bound on $\|\Delta A_{j}^{(i)}\|_{\mathrm{F}}$ holds uniformly for all $i$, leading to $\|F_{i+1,i}-\mathbf{F}_{i+1,i}\|_{\mathrm{F}}\leq C_{1}\,\Delta t^{(i)},$ for some constant $C_{1}$ independent of $i$.
The cumulative deviation over $l$ steps can be written as
$F_{i+l,i}-\mathbf{F}_{i+l,i}=\sum_{m=1}^{l}\mathbf{F}_{i+l,i+m}\,\big(F_{i+m,i+m-1}-\mathbf{F}_{i+m,i+m-1}\big)\,F_{i+m-1,i}.$
Taking norms and applying the triangle inequality yields
$\|F_{i+l,i}\|_{\mathrm{F}}$ $\leq\|\mathbf{F}_{i+l,i}\|_{\mathrm{F}}+\|\mathbf{F}_{i+l,i+m}\|_{\mathrm{F}}C_{1}\sum_{m=1}^{l}\Delta t^{(i+m-1)}\|F_{i+m-1,i}\|_{\mathrm{F}}$
$\leq C_{0}+C_{0}C_{1}\sum_{m=1}^{l}\Delta t^{(i+m-1)}\|F_{i+m-1,i}\|_{\mathrm{F}}.$
Finally, using the discrete Gronwall–Bellman inequality [22], we obtain
$\|F_{i+l,i}\|_{\mathrm{F}}\leq C_{0}e^{C_{0}C_{1}T}\equiv C_{F}.$
Thus the discrete STM remains uniformly bounded within each period. ∎
Lemma 5.
Under assumptions A1-A3, let $\Psi(t_{p},t_{0})$ and $F_{p,0}$ denote the auxiliary and discrete STM over $t_{0},t_{p}]$. Consider the difference $\Psi(t_{p},t_{0})-F_{p,0}$ acting on the normalized eigenvector matrices $\mathcal{U}_{\cdot}(t_{0})$ of $\Psi(t_{p},t_{0})$, with $\lVert\mathcal{U}_{\cdot}(t_{0})\rVert_{\mathrm{F}}=1$. Then the following bounds hold:
  1. 1.
    $\left\lVert\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{1}(t_{0})\right\rVert_{\mathrm{F}}\leq C_{1}\Delta t^{s}$,
  2. 2.
    $\left\lVert\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{2}(t_{0})\right\rVert_{\mathrm{F}}\leq C_{2}\Delta t$.
with constants $C_{1}$, $C_{2}$ independent of $\Delta t$.
Proof.
We decompose the global difference into sums of local difference:
$\left[\Psi(t_{p},t_{0})-F_{p,0}\right]\mathcal{U}_{\tau}(t_{0})$ (41)
$=$ $\sum_{i=1}^{p}F_{p,i}\cdot\left[\Psi(t_{i},t_{i-1})-F_{i,i-1}\right]\cdot\Psi_{i-1,0}\cdot\mathcal{U}_{\tau}(t_{0})$
[…]

### v2

@@ 4.4 Convergence Analysis
$\operatorname{sep}_{2}(A_{11},A_{22})=\mathrm{d}(A_{11},A_{12})\equiv\min\{\left\lvert\lambda-\mu\right\rvert,\,\lambda\in\lambda(A_{11}),\mu\in\lambda(A_{22})\}.$
With the global difference estimate and the invariant subspace perturbation theorem at hand, we are now in a position to complete the proof of Theorem˜4.1.
Proof 4.8.
The three statements are proved by the same perturbation argument based on the block representation of $\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})$ and a spectral separation estimate. We therefore present the proof for part (3), which contains the main ideas. Parts (2) and (1) follow by the same reasoning with minor simplifications, as outlined at the end of the proof.
Let $1\leq k\leq n$, we partition $\mathcal{X}(t_{0})$ as $\mathcal{X}(t_{0})=\begin{bmatrix}X_{k}(t_{0})&X_{n-k}(t_{0})\end{bmatrix}$ After the similarity transformation induced by the moving basis $\mathcal{Z}(t_{0})$, the one-period stacked propagator admits the block representation Eq.˜29:
$\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})$
$=$ $\begin{bmatrix}e^{M_{k}T}&&\\ &e^{M_{n-k}T}\\ &&D^{p}\end{bmatrix}+\begin{bmatrix}e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{n-k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{Y}}\\ e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{n-k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{Y}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{k}}&D^{p}E_{\mathcal{Y}\mathcal{X}_{n-k}}&D^{p}E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}$.
For brevity, we set
$E_{11}=e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}},\quad E_{12}=\begin{bmatrix}e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{n-k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{Y}}\end{bmatrix}$,
$E_{21}=\begin{bmatrix}e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{k}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{k}}\end{bmatrix},\quad E_{22}=\begin{bmatrix}e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{n-k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{Y}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{n-k}}&D^{p}E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}$.
Since all Floquet multipliers $e^{\mu_{k}T}$ are assumed semisimple and sorted by non-increasing order, thus $e^{\mu_{n}T}$ is the smallest magnitude Floquet multiplier. Since $\|D^{p}\|=\nu_{\max}^{p}\to 0$ as $p\to\infty$, there exists $p_{0}\in\mathbb{N}$ such that $\|D^{p}\|\leq e^{\mu_{n}T}/2$ for all $p\geq p_{0}$. Because both $e^{MT}$ and $D^{p}$ are diagonal (hence normal), their spectra are separated, and
$\operatorname{sep}(e^{MT},D^{p})=\operatorname{dist}\!\bigl(\lambda(e^{MT}),\lambda(D^{p})\bigr)\geq e^{\mu_{n}T}/2,\qquad p\geq p_{0}.$
For sufficient large $p$, we have
$\begin{aligned} \delta=&\operatorname{sep}(e^{M_{k}T},\begin{bmatrix}e^{M_{n-k}}&\\ &D^{p}\end{bmatrix})-\left(\left\lVert e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}}\right\rVert+\left\lVert\begin{bmatrix}E_{11}&\\ &E_{22}\end{bmatrix}\right\rVert\right)\\ >&1/2\left\lvert e^{\mu_{m}T}-e^{\mu_{m+1}T}\right\rvert\end{aligned}$,
and
$\left\lVert E_{21}\right\rVert\left(\left\lVert 0\right\rVert+\left\lVert E_{21}+E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}.$
Thus, the conditions of Theorem˜4.6 are satisfied, and there exists a matrix $R\in\mathbb{C}^{((d-1)n+k)\times k}$ with
$\|R\|\leq C\,\nu_{\max}^{p},$
such that
$\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})\begin{bmatrix}I_{k}\\ R\end{bmatrix}=\begin{bmatrix}I_{k}\\ R\end{bmatrix}B,$
where the reduced $n\times n$ matrix $B$ is given by
[…]